OpenStudy (anonymous):

Is there any way to write 2/3ln(x-1/x=1) as ln(x-1/x+1)^3/2? My answer for an integral problem was the 1st mentioned answer and trying to see if I just missed log properties of writing the answer a different way.

hartnn (hartnn):

\(\huge a\log b = \log b^a\)

hartnn (hartnn):

so you should have got 3/2 ln (x-1/x+1)

OpenStudy (anonymous):

Maybe I had an error in my calculations. Can you check a partial fraction integral with me if I show you what I did?

hartnn (hartnn):

ofcourse :)

hartnn (hartnn):

post original Question also

OpenStudy (anonymous):

\[\int\limits 3/x^2-1 dx\] \[3/(x-1)(x+1)=A/x+1 + B/x-1\]

OpenStudy (anonymous):

can I skip ahead to my partial fractions?

hartnn (hartnn):

yes, what u got as A , B ?

OpenStudy (anonymous):

A= -2/3 B=2/3

hartnn (hartnn):

correct :)

OpenStudy (anonymous):

So... \[\int\limits -2/3/(x+1) dx + \int\limits 2/3 (x-1) dx\]

hartnn (hartnn):

oh wait!

hartnn (hartnn):

A= -3/2 B = 3/2

hartnn (hartnn):

3 = A(x-1) +B(x+1) right ?

OpenStudy (anonymous):

yes

hartnn (hartnn):

if you put x=1 here, u get 3 = B (2) B=3/2 :)

OpenStudy (anonymous):

let me check that out (;

OpenStudy (anonymous):

Awesome!! Thank you so much for your help (; What school do you go to?

hartnn (hartnn):

i completed my education this year :)

OpenStudy (anonymous):

In math?

hartnn (hartnn):

nopes, in Telecommunication system Engineering, but it had good amount of math too :)

OpenStudy (anonymous):

Have you found a good job since getting out of school? Im working on EE

hartnn (hartnn):

yes, i did! i work in Cisco Systems as software engineer. but i live in India, the situation in your country might be completely different...

OpenStudy (anonymous):

Thats great! I live in U.S. One last question about partial fractions...m just starting ot do these problems and have been wondering if it matters which factor such as (x-1) (x+1) I use as denominator of A and B?

hartnn (hartnn):

no it does not matter. you will just get different (exchanged) values of A,B but your final answer will be same :)

OpenStudy (anonymous):

Thank you so much I really appreciate your help. Do you come on here to practice your math skills?

hartnn (hartnn):

you're welcome ^_^ i come here mostly to help people with their math, i like helping, and math :)

OpenStudy (anonymous):

Thanks again (;